A group of 8 students in your class having the weight of 50, 52, 53, 54, 56, 57, 58, and 60.

1.Calculate the standard deviation of the sampling distribution of the sample means. Compare this to the standard deviation of the population. 


1
Expert's answer
2022-05-26T14:00:35-0400

standard deviation of the sampling distribution of the sample means:

"S=\sqrt{\frac{\sum(x_i-\=x)^2}{n-1}}"

"\=x=\frac{50+52+53+54+56+57+58+60}8=55"

"S=\sqrt{\frac{(50-55)^2+(52-55)^2+(53-55)^2+(54-55)^2+(56-55)^2+(57-55)^2+(58-55)^2+(60-55)^2}{7}}=\sqrt{11.143}=3.34"

standard deviation of the population:

"\sigma=\sqrt{\frac{\sum(x_i-\mu)^2}{N}}\\=\sqrt{\frac{78}{8}}=3.12"

As we can see, the standard deviation of the population is less than the standard deviation of the sampling distribution of the sample means due to the differences in formulas: in standard deviation of the population we divide by N (count of values in population), while in standard deviation of the sampling distribution of the sample means we divide by n-1 (count of values in sample-1).